OpenStudy (anonymous):

(2x^2-3x/2x-10)*(4/3(2x-3)

TheSmartOne (thesmartone):

\[\frac{ 2x^2-3x }{ 2x-10} \times \frac{ 4 }{ 3 }(2x-3)\]

TheSmartOne (thesmartone):

What are you supposed to do? @fernie2020

OpenStudy (triciaal):

2x^2 -3x = x(2x - 3 ) 2x-10 = 2(x-5) rewrite the original using the factors then simplify hint some factors will cancel

OpenStudy (anonymous):

so after i did that, I was left with  @triciaal